OpenStudy (anonymous):

The senior class at high school A and high school B planned separate trips to New York City. The senior class at high school A rented and filled 1 van and 6 buses with 372 students. High school B rented and filled 4 vans and 12 buses with 780 students. Each van and each bus carried the same number of students. How many students can a van carry? How many students can a bus carry?

OpenStudy (lilg132):

372 = V + 6B 780 = 4V + 12B 372 - 6B = V 780 = 372 - 6B + 12B 780 - 372 = -6B + 12B 408 = 6B 408 / 6 = 6B / 6 68 = B ONE BUS CARRIES 68 PEOPLE NOW TO WORK OUT HOW MUCH ONE VAN DOES 372 = V + 6B 780 = 4V + 12B 372 - V = 6B 780 = 4V + 372 - V 780 - 372 = 4V - V 408 = 3V 408 / 3 = 3V / 3 136 = V ONE VAN CARRIES 136 PEOPLE V = VAN B = BUS dont forget to click best response =)